OpenStudy (anonymous):

Which equation is not equivalent to the formula r = st? A) s = r/t B) t = r/s C) s = t/r D) r = ts

OpenStudy (lstrasmann):

Well, the commutative property would make D be equivalent, so we can cross that out.

OpenStudy (anonymous):

Which ones can we cross out?

OpenStudy (lstrasmann):

B can also be crossed out. \[r=st\]\[\frac{ r }{ s }=\frac{ st }{ s }\]\[\frac{ r }{ s }=t\]